OpenStudy (anonymous):

find all solutions of the equation on the interval [0,2pie) cos(x+pie/4)-cos(x-pie/4)=1

OpenStudy (anonymous):

\[\cos(x+\frac{\pi}{4})-\cos(x-\frac{\pi}{4})=1\]multiply by \(\frac{\sqrt{2}}{2}\)